Angelina Rogers was looking at different simple discount notes available to her. She found a note that would cost her $135 in interest at 8% for 60 days. What was the face value?
Answer by rfer(16322)

135=P*0.08*60/365
135/(0.08*60/365)=P
135/0.01315=P
$10266.16=P